# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v4

@@ 2.2 Lattice Problems
The $\mathsf{M\text{-}SIS}_{n,k,\ell,q,\beta}$ problem is: given a uniformly random matrix $\mathbf{A}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k\times\ell}$, find a non-zero vector $\mathbf{z}\in R_{q}^{\ell}$ such that $\mathbf{A}\mathbf{z}=\mathbf{0}\mod q$ and $\left\|\mathbf{z}\right\|_{\infty}\leq\beta$.
@@ 2.3 ML-DSA (FIPS 204)
The z-bound check and r0-check together yield $\approx 20$$25\%$ per-attempt success (4–5 expected attempts). This is inherent to Fiat-Shamir with Aborts [Lyu12], not threshold-specific.
@@ 2.4 Shamir Secret Sharing
Shamir’s $(T,N)$-threshold secret sharing [Shamir79] allows a secret $s\in\mathbb{Z}_{q}$ to be distributed among $N$ parties such that any $T$ parties can reconstruct $s$, but any T1T-1 parties learn nothing.
@@
$\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\mod q$
@@
For the standard evaluation points $\{1,2,\ldots,T\}$, Lagrange coefficients satisfy $\lambda_{i}=(-1)^{i-1}\binom{T}{i}$, growing as $\approx 2^{T}/\sqrt{T}$. For $T\leq 8$, coefficients remain small ($\binom{8}{4}=70$), enabling short-coefficient schemes [dPN25]. For $T\geq 24$, coefficients exceed $q\approx 2^{23}$, aligning with the Ball-Çakan-Malkin bound [BCM21]. The obstacle is that $\lambda_{i}\cdot\mathbf{z}_{i}$ has norm $O(q\cdot\gamma_{1})\gg\gamma_{1}-\beta$, causing rejection.
@@ A $(T,N)$-threshold signature scheme consists of algorithms:
With Shamir nonce DKG and coordinator-based communication (Remark 3.7), Profiles P1 and P3+ require only $|S|\geq T$ signers—the information-theoretic minimum for Shamir reconstruction. This improves upon standard pairwise-mask schemes [CGGMP20, DKOSS24] that require $|S|\geq T+1$. The nonce DKG provides statistical share privacy (no computational assumptions) even with a single honest party (see Theorem 4.4 and Remark 4.1). In P2’s broadcast model, $|S\setminus C|\geq 2$ is required for commitment privacy (Lemma 4.9).
@@ 3.1 Overview
The core challenge in threshold ML-DSA is handling the large Lagrange coefficients that arise in Shamir secret sharing over $\mathbb{Z}_{q}$. As noted in Remark 2.3, the Lagrange coefficients grow exponentially as $\approx 2^{T}$, exceeding $q$ for $T\geq\log_{2}q\approx 23$. Even for moderate $T$ (e.g., $T=16$), the coefficients $\binom{16}{8}=12870$ can cause the naive approach to inflate norms beyond the rejection threshold.
@@ 3.3 Setup Phase
Before key generation, each pair of parties $(i,j)$ with $i<j$ establishes a shared secret seed $\mathsf{seed}_{i,j}\in\{0,1\}^{256}$. This can be done via:
  • ML-KEM key exchange: Each party generates an ML-KEM keypair, and pairs perform encapsulation/decapsulation to derive shared secrets.
  • Pre-shared keys: In some settings, seeds may be distributed by a trusted setup.
@@
5:Compute $\mathbf{t}_{1}=\mathsf{HighBits}(\mathbf{t},2^{d})$
@@
7:for each polynomial $f$ in $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ do
8:  Create Shamir sharing: $f(X)=f+r_{1}X+\cdots+r_{T-1}X^{T-1}$
9:  Compute shares: $f_{i}=f(i)$ for $i\in[N]$
@@ 3.4 Shamir Nonce DKG
The coefficients $\mathbf{a}_{i,k}$ for $k\geq 1$ are sampled uniformly over $R_{q}^{\ell}$ (not from a short distribution). This is essential: the adversary observes T1T-1 evaluations of $\mathbf{f}_{h}$ (the honest party’s polynomial), leaving exactly one degree of freedom. The uniform prior distribution of $\mathbf{a}_{h,T-1}$ ensures high entropy in this free parameter; combined with the bounded-nonce constraint, the honest party’s share $\mathbf{f}_{h}(x_{h})$ retains conditional min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits per session. See Theorem 4.4 and Remark 4.6.
@@
The nonce DKG requires $|S|\cdot(|S|-1)$ point-to-point messages, each of size $n\ell\cdot\lceil\log_{2}q\rceil$ bits ($\approx 3.8$ KB for ML-DSA-65). For a $(3,5)$-threshold with $|S|=3$, this is 6 messages totaling $\approx 23$ KB. The DKG round is independent of the message to be signed and can be executed offline as a preprocessing step, so online signing latency is unchanged.
@@ 3.5 Pairwise-Canceling Masks
For signing set $S$ and domain separator $\mathsf{dom}\in\{0,1\}^{*}$, party $i$ computes:
$\mathbf{m}_{i}^{(\mathsf{dom})}=\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{dom})$
where $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{d}$ is instantiated with SHAKE-256, and $d$ is the appropriate dimension ($k$ for commitment masks, $k$ for r0-check masks).
We use domain separators $\mathsf{dom}=\mathsf{nonce}\|\texttt{"comm"}$ for commitment masks $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$ and $\mathsf{dom}=\mathsf{nonce}\|\texttt{"s2"}$ for r0-check masks $\mathbf{m}_{i}^{(s2)}\in R_{q}^{k}$.
@@
Proof 3.5.
Each pair $(i,j)$ with $i<j$ contributes:
  • $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})$ from party $i$ (since $j>i$)
  • $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})$ from party $j$ (since $i<j$)
These cancel exactly, so $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
@@ 3.6 Signing Protocol
19:  Compute Lagrange coefficient $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\bmod q$
@@
24: Aggregate $\mathbf{W}=\sum_{j\in S}\mathbf{W}_{j}=\sum_{j}\lambda_{j}\mathbf{w}_{j}+\sum_{j}\mathbf{m}_{j}^{(w)}=\mathbf{A}\mathbf{y}=\mathbf{w}$ $\triangleright$ Masks cancel
25: Compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$
26: Compute $\tilde{c}=H_{\mathsf{chal}}(\mu\|\mathbf{w}_{1})$ and expand to challenge $c\in R_{q}$ $\triangleright$ FIPS 204 spec
27:if $c$ is not invertible in $R_{q}$ then
28:  return $\bot$ (challenge invertibility abort, retry from Round 1) $\triangleright$ Negligible probability
29:end if
@@
45:r0-check: Verify $\left\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\right\|_{\infty}<\gamma_{2}-\beta$
46:if r0-check fails then
47:  return $\bot$ (r0 abort, retry from Round 1)
48:end if
49:Compute $\mathbf{r}=\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d}$ $\triangleright$ $\mathbf{w}-c\mathbf{s}_{2}+c\mathbf{t}_{0}$
50:Compute $c\mathbf{t}_{0}$ using $c\mathbf{s}_{2}$ from r0-check $\triangleright$ See Subsection 3.10
51:Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{r},2\gamma_{2})$
52:if $\mathsf{weight}(\mathbf{h})>\omega$ then
53:  return $\bot$ (hint weight abort)
54:end if
55:return signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
@@
Remark 3.9 (Nonce Range Rounding).
When $\gamma_{1}$ is not divisible by $|S|$, rounding slightly reduces the nonce range. This has negligible impact: for ML-DSA-65 with $|S|=17$, the range loss is $2\times 10^{-5}$ per coordinate and does not affect the z-bound check since $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor\geq\gamma_{1}-(|S|-1)>\gamma_{1}-\beta$ (holds for all $|S|\leq\beta+1=197$, covering all practical configurations). See Appendix G for details.

### v5

@@ 2.2 Lattice Problems
The $\mathsf{M\text{-}SIS}_{n,k,\ell,q,\beta}$ problem is: given a uniformly random matrix $\mathbf{A}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k\times\ell}$, find a non-zero vector $\mathbf{z}\in R_{q}^{\ell}$ such that $\mathbf{A}\mathbf{z}=\mathbf{0}\mod q$ and $\left\|\mathbf{z}\right\|_{\infty}\leq\beta$, where $\left\|\cdot\right\|_{\infty}$ denotes the $\ell^{\infty}$ norm computed with polynomial coefficients centered in (q/2,q/2](-q/2,q/2].
Remark 2.2 (Notation: $\beta$ disambiguation).
The parameter $\beta$ in the M-SIS definition above is a generic norm bound following standard lattice notation [LPR10]. The ML-DSA signing parameter $\beta=\tau\eta=196$ (see parameters below) is a distinct, concrete value: the SIS norm bound instantiated for our unforgeability proof is $\gamma_{1}-\beta_{\mathsf{sign}}=2^{19}-196\approx 524{,}092$ (Remark 2.3). No clash arises because the SIS definition is parameterized generically and the ML-DSA $\beta=196$ is always used in the signing context.
Remark 2.3 (SelfTargetMSIS).
The security reduction in Theorem 4.2 extracts a solution to the inhomogeneous SelfTargetMSIS variant [KLS18]: given $\mathbf{A}|\!-\!\mathbf{t}_{1}\cdot 2^{d}]$, find a short $(\mathbf{v},\Delta c)$ with $\|\mathbf{v}\|_{\infty}<\gamma_{1}-\beta$ such that $\mathbf{A}|\!-\!\mathbf{t}_{1}\cdot 2^{d}]\bigl[\begin{smallmatrix}\mathbf{v}\\ \Delta c\end{smallmatrix}\bigr]=2\gamma_{2}\mathbf{w}_{1}+\mathbf{r}_{0}$ where $\mathbf{w}_{1}$ is a known target and $\|\mathbf{r}_{0}\|_{\infty}\leq\gamma_{2}$ (the $\mathsf{LowBits}$ residual from $\mathsf{HighBits}$ rounding). By [KLS18], SelfTargetMSIS hardness follows from M-SIS hardness under a standard parameter embedding; we therefore state Theorem 4.2 in terms of M-SIS.
@@ 2.3 ML-DSA (FIPS 204)
The z-bound check and r0-check together yield $\approx 20$$25\%$ per-attempt success (4–5 expected attempts). This is inherent to Fiat-Shamir with Aborts [Lyu12], not threshold-specific. Note: In our threshold protocol, the aggregated nonce follows an Irwin-Hall distribution (sum of $|S|$ bounded uniform variables) rather than a single uniform draw. The Irwin-Hall distribution has greater density near zero, shifting more nonces into the acceptance region $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$. As a result, observed threshold signing success rates of 21–45% are consistent with (and can slightly exceed) the single-signer baseline; see Table 5.
@@ 2.4 Shamir Secret Sharing
Shamir’s $(T,N)$-threshold secret sharing [Shamir79] allows a secret $s\in\mathbb{Z}_{q}$ to be distributed among $N$ parties such that any $T$ parties can reconstruct $s$, but any T1T-1 parties learn nothing. For vector (module) secrets $\mathbf{s}\in R_{q}^{\ell}$, sharing is applied coordinate-by-coordinate over $\mathbb{Z}_{q}$.
@@
$\lambda_{i}=\prod_{j\in S,j\neq i}j\cdot(j-i)^{-1}\bmod q$
@@
For the standard evaluation points $\{1,2,\ldots,T\}$, Lagrange coefficients satisfy $\lambda_{i}=(-1)^{i-1}\binom{T}{i}$, growing as $\approx 2^{T}/\sqrt{\pi T/2}$ (by Stirling’s formula). For $T\leq 8$, coefficients remain small ($\binom{8}{4}=70$), enabling short-coefficient schemes [dPN25]. For $T\geq 26$, coefficients exceed q=8,380,417q=8{,}380{,}417 (since $\binom{25}{12}=5{,}200{,}300<q$ but $\binom{26}{13}=10{,}400{,}600>q$), aligning with the Ball-Çakan-Malkin bound [BCM21]. For short-coefficient LSSS designs [dPN25, BCM21] that require Lagrange products $\lambda_{i}\cdot\mathbf{z}_{i}$ to remain short before modular reduction, this is a binding obstacle: the product has norm $O(q\cdot\gamma_{1})\gg\gamma_{1}-\beta$. When $|S|>T$, the maximum Lagrange coefficient magnitude grows as $\binom{|S|}{\lfloor|S|/2\rfloor}$ (larger than for $|S|=T$), so the constraint on $T$ is the binding one in practice. In our construction, shares $\mathbf{z}_{i}\in\mathbb{Z}_{q}^{\ell}$ are large, but the aggregate satisfies $\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$ by the Shamir reconstruction identity; since $\|\mathbf{y}+c\mathbf{s}_{1}\|_{\infty}\leq\gamma_{1}-\beta<q/2$, the computation in $\mathbb{Z}_{q}$ yields the same result as integer arithmetic, and the z-bound check operates on the correct short value.
@@ A $(T,N)$-threshold signature scheme consists of algorithms:
With Shamir nonce DKG and coordinator-based communication (Remark 3.9), Profiles P1 and P3+ require only $|S|\geq T$ signers—the minimum required by Shamir’s $(T,N)$ threshold property for reconstruction. This improves upon standard pairwise-mask schemes [CGGMP20, DKOSS24] that require $|S|\geq T+1$. The nonce DKG provides statistical share privacy (no computational assumptions) even with a single honest party (see Theorem 4.7 and Remark 4.1). In P2’s broadcast model, $|S\setminus C|\geq 2$ is required for commitment privacy (Lemma 4.15).
@@
Remark 2.10 (Informal vs. UC Privacy).
Definition 2.9 is an informal stand-in. The full UC-level privacy guarantees, with explicit simulators and hybrid arguments, are established separately for each deployment profile: Theorem 4.21 (P1), Theorem 6.3 (P2), and Theorem 6.6 (P3+).
@@ 3.1 Overview
The core challenge in threshold ML-DSA is handling the large Lagrange coefficients that arise in Shamir secret sharing over $\mathbb{Z}_{q}$. As noted in Remark 2.5, the Lagrange coefficients for standard evaluation points $\{1,\ldots,N\}$ can be as large as $\binom{T}{\lfloor T/2\rfloor}\approx 2^{T}/\sqrt{T}$ over $\mathbb{Z}$ before modular reduction—for example, $\binom{16}{8}=12870$ at $T=16$ (see Lemma Q.30). When the naive reconstruction $\mathbf{z}=\sum_{i\in S}\lambda_{i}\mathbf{z}_{i}$ is computed, each response share $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ is scaled by $\lambda_{i}$, inflating the $\ell_{\infty}$-norm by up to $12870\times$ (at T=16T=16)—far beyond the rejection-sampling bound $\gamma_{1}-\beta=524092$. For large $T$, the coefficient magnitude eventually exceeds $q$, causing aliasing in the modular reduction.
@@ 3.2 Threat Model and Deployment Profiles
(a)  P1 — TEE-Assisted
[…]
(b)  P2 — Fully Distributed
[…]
(c)  P3+ — Semi-Async 2PC
[…]
Figure 1: Protocol message flows for the three deployment profiles. Shaded top regions are offline preprocessing; dashed lines are actor lifelines. (a) P1: masked commitments $\hat{\mathbf{W}}_{i}$ sent only to trusted TEE coordinator, enabling $|S\setminus C|=1$. (b) P2: all values broadcast; dishonest-majority security via SPDZ+edaBits; r0-check computed inside MPC. (c) P3+: semi-async signers respond within window $\Delta$; two CPs run garbled-circuit 2PC for r0-check. LR = Logical Round; R = Protocol Round.
@@ 3.3 Setup Phase
Before key generation, each pair of parties $(i,j)$ with $i<j$ establishes a shared secret seed $\mathsf{seed}_{i,j}\in\{0,1\}^{256}$. This may be done via ML-KEM key exchange (each party generates a keypair; pairs perform encapsulation and decapsulation to derive the shared seed) or via pre-shared keys from a trusted setup.
@@
Remark 3.1 (Key Generation Scope).
Threshold key generation is an application-layer component outside the scope of FIPS 204. The resulting public key $\mathsf{pk}=(\rho,\mathbf{t}_{1})$ is byte-identical to a standard FIPS 204 ML-DSA-65 public key and is usable with any FIPS 204-compliant verification implementation without modification. The threshold secret key shares $\{\mathsf{sk}_{i}\}$ are specific to this construction and replace the single-party FIPS 204 signing key; they do not affect the signature format or verifiability.
@@
5:Decompose $(\mathbf{t}_{1},\mathbf{t}_{0})\leftarrow\mathsf{Power2Round}(\mathbf{t},d)$ $\triangleright$ FIPS 204 Alg. 29; $\mathbf{t}=\mathbf{t}_{1}\cdot 2^{d}+\mathbf{t}_{0}$
@@
7:for each secret polynomial $s$ in $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ do
8:  Create Shamir sharing: $p_{s}(X)=s+a_{1}X+\cdots+a_{T-1}X^{T-1}$,  $a_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}$
9:  Compute shares: $s_{i}=p_{s}(i)$ for $i\in[N]$
@@ 3.4 Shamir Nonce DKG
The coefficients $\mathbf{a}_{i,k}$ for $k\geq 1$ are sampled uniformly over $R_{q}^{\ell}$ (not from a short distribution). This is essential: the adversary observes T1T-1 evaluations of $\mathbf{f}_{h}$ (the honest party’s polynomial), leaving exactly one degree of freedom. The uniform prior distribution of $\mathbf{a}_{h,T-1}$ ensures high entropy in this free parameter; combined with the bounded-nonce constraint, the honest party’s share $\mathbf{f}_{h}(x_{h})$ retains conditional min-entropy $\geq n\ell\cdot\log_{2}\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$ bits per session (the support of each per-party share coordinate is $\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}$, a range of $2\lfloor\gamma_{1}/|S|\rfloor+1$ values counting both endpoints; the approximation $\approx n\ell\cdot(20-\log_{2}|S|)$ bits holds to within $1$ bit for $|S|\leq 17$). See Theorem 4.7 and Remark 4.10.
@@
The nonce DKG requires $|S|\cdot(|S|-1)$ point-to-point messages, each of size $n\ell\cdot\lceil\log_{2}q\rceil$ bits ($\approx 3.6$ KB for ML-DSA-65; $256\times 5\times 23=29{,}440$ bits). For a $(3,5)$-threshold with $|S|=3$, this is 6 messages totaling $\approx 22$ KB; for $|S|=32$, the DKG requires $32\times 31=992$ messages totaling $\approx 3.6$ MB. This $O(|S|^{2})$ offline preprocessing cost does not appear in the online round complexity comparison; it is amortizable over batches of signing sessions (multiple nonce DKG rounds can be preprocessed during idle time) and is absent from the online communication cost ($\approx 12.3$ KB per party per online attempt). The DKG round is independent of the message to be signed and can be executed offline as a preprocessing step, so online signing latency is unchanged.
Remark 3.4 (Why No VSS for the Nonce DKG).
The key DKG uses Feldman commitments [Feldman87] because the long-term secret $\mathbf{s}_{1}$ has short coefficients and parties must verifiably bind their contributions during a one-time setup. The nonce DKG has a different structure: the higher-degree polynomial coefficients are sampled $R_{q}^{\ell}$-uniformly (by design, for entropy), and the constant term $\hat{\mathbf{y}}_{i}$ must remain hidden to preserve nonce privacy (Theorem 4.7). Feldman-style commitments publish $g^{a_{k}}$ for each coefficient, which would not hide $\hat{\mathbf{y}}_{i}$ against a discrete-log adversary and would require a hiding commitment scheme instead—substantially complicating the offline round. We instead adopt an optimistic approach: shares are sent over private authenticated channels, and any cheating party is detectable a posteriori by the blame protocol (Section 6.1), which asks parties to reveal their full polynomials to the TEE. The TEE recomputes each share and verifies consistency with the submitted $\mathbf{z}_{i}$ values. The optimistic threshold of $K=33$ consecutive aborts ensures blame is triggered before a malicious party can cause meaningful disruption (honest probability $(0.75)^{33}<10^{-4}$).
@@ 3.5 Pairwise-Canceling Masks
For signing set $S$ and PRF context $\mathsf{ctx}\in\{0,1\}^{*}$ (composed as $\mathsf{ctx}=\mathsf{nonce}^{*}\|\mathsf{label}$, where $\mathsf{nonce}^{*}$ is the appropriate per-session identifier defined below, and $\mathsf{label}$ is a fixed domain tag), party $i$ computes:
$\mathbf{m}_{i}^{(\mathsf{ctx})}=\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{ctx})$
where $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{k}$ is instantiated with SHAKE-256 (here $k=6$ is the module rank, not the dropping-bits parameter $d=13$). The per-session nonce ensures masks are fresh in each signing session (see Mask Hiding, Lemma 4.15).
We use $\mathsf{ctx}=\mathsf{nonce}_{0}\|\texttt{"comm"}$ for commitment masks $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$, where $\mathsf{nonce}_{0}=H(\texttt{"nonce0"}\|\mathsf{Com}_{i_{1}}\|\cdots)$ is the pre-challenge nonce derived at the start of Round 2 (before the challenge $c$ is known). We use $\mathsf{ctx}=\mathsf{nonce}\|\texttt{"s2"}$ for r0-check masks $\mathbf{m}_{i}^{(s2)}\in R_{q}^{k}$, where $\mathsf{nonce}$ is the post-challenge per-session identifier.
@@
Proof 3.7.
Each pair $(i,j)$ with $i<j$ contributes $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})$ from party $i$ (since $j>i$) and $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})$ from party $j$ (since $i<j$); these cancel exactly in the sum, so $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
@@ 3.6 Signing Protocol
19:  Compute Lagrange coefficient for reconstruction at point 0: $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\bmod q$ $\triangleright$ Shamir interpolation at $x=0$
@@
24: Aggregate $\mathbf{W}=\sum_{j\in S}\mathbf{W}_{j}=\sum_{j}(\lambda_{j}\mathbf{w}_{j}+\mathbf{m}_{j}^{(w)})$ $\triangleright$ Substitute definition of $\mathbf{W}_{j}$
25:  $\sum_{j}\lambda_{j}\mathbf{w}_{j}+\sum_{j}\mathbf{m}_{j}^{(w)}=\sum_{j}\lambda_{j}\mathbf{A}\mathbf{y}_{j}+\mathbf{0} \triangleright$ Masks cancel (Lemma 3.6); $\mathbf{w}_{j}=\mathbf{A}\mathbf{y}_{j}$
26:  $\mathbf{A}\sum_{j}\lambda_{j}\mathbf{y}_{j}=\mathbf{A}\mathbf{y} \triangleright$ Linearity of $\mathbf{A}$; Lagrange reconstruction of $\mathbf{y}$ from $\{\mathbf{y}_{j}\}_{j\in S}$
27: Set $\mathbf{w}\leftarrow\mathbf{W}$ $\triangleright$ $\mathbf{w}=\mathbf{A}\mathbf{y}$; used in r0-check and hint computation
28: Compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{w},2\gamma_{2})$
29: Compute $\tilde{c}=H_{\mathsf{chal}}(\mu\|\mathbf{w}_{1})$ and expand to challenge $c\in R_{q}$ $\triangleright$ FIPS 204 spec; no invertibility check needed
@@
45:Combiner computes $c\mathbf{s}_{2}\leftarrow\sum_{i\in S}\mathbf{V}_{i}=\sum_{i\in S}(\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)})$ $\triangleright$ Substitute definition of $\mathbf{V}_{i}$
46:  $c\sum_{i\in S}\lambda_{i}\mathbf{s}_{2,i}+\sum_{i\in S}\mathbf{m}_{i}^{(s2)}=c\mathbf{s}_{2}+\mathbf{0} \triangleright$ Lagrange reconstruction of $\mathbf{s}_{2}$ from shares $\{\mathbf{s}_{2,i}\}_{i\in S}$ using same $\lambda_{i}$ as Line 206; masks cancel (Lemma 3.6)
47:r0-check: Verify $\left\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\right\|_{\infty}<\gamma_{2}-\beta$
48:if r0-check fails then
49:  return $\bot$ (r0 abort, retry from Round 1)
50:end if
51:Compute $\mathbf{r}=\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d}$ $\triangleright$ $\mathbf{w}-c\mathbf{s}_{2}+c\mathbf{t}_{0}$
52:Set $c\mathbf{t}_{0}\leftarrow\mathbf{w}-\mathbf{r}-c\mathbf{s}_{2} \triangleright$ $\mathbf{w}-(\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})-c\mathbf{s}_{2}$
53:Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{r},2\gamma_{2})$
54:if $\mathsf{weight}(\mathbf{h})>\omega$ then
55:  return $\bot$ (hint weight abort)
56:end if
57:return signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
@@
Individual masked commitments $\mathbf{W}_{i}=\lambda_{i}\mathbf{A}\mathbf{y}_{i}+\mathbf{m}_{i}^{(w)}$ are sent only to the coordinator, not broadcast to all parties. This is critical for privacy when $|S\setminus C|=1$: if $\mathbf{W}_{h}$ were broadcast and the adversary controls all other parties in $S$, the pairwise mask $\mathbf{m}_{h}^{(w)}$ would be fully computable (all seeds are shared with corrupted parties), revealing $\lambda_{h}\mathbf{A}\mathbf{y}_{h}$.
Recovery of $\mathbf{y}_{h}$ from $\mathbf{A}\mathbf{y}_{h}$ (when $k>\ell$). For ML-DSA-65, the public matrix $\mathbf{A}\in R_{q}^{k\times\ell}$ with $k=6$, $\ell=5$ (so $k>\ell$) is generated via the random oracle ExpandA. The Module-LWE assumption [LS15, ADPS16] requires that $\mathbf{A}$ has full column rank $\ell$ with overwhelming probability over the random oracle sampling (see [FIPS204, Section 3.4] for ExpandA). When $\mathbf{A}$ has full column rank and $k\geq\ell$, the left inverse $((\mathbf{A}^{\top}\mathbf{A})^{-1}\mathbf{A}^{\top})$ exists over $R_{q}$ (modulo invertibility of $\mathbf{A}^{\top}\mathbf{A}$, which holds with overwhelming probability for random $\mathbf{A}$), enabling recovery:
$\mathbf{y}_{h}=(\mathbf{A}^{\top}\mathbf{A})^{-1}\mathbf{A}^{\top}\cdot(\lambda_{h}\mathbf{A}\mathbf{y}_{h})/\lambda_{h}.$
The scalar $\lambda_{h}$ is invertible modulo $q$ (as all Lagrange coefficients are, by construction). Thus, revealing $\mathbf{A}\mathbf{y}_{h}$ enables key extraction via $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ (the adversary knows $\mathbf{z}_{h}$ from the public signature and $c$ from the challenge). This is a non-trivial lattice attack that does not rely on the high min-entropy of $\mathbf{y}_{h}$—it is a direct algebraic recovery, bypassing statistical privacy. Therefore, $\mathbf{W}_{i}$ must remain coordinator-only.
@@
Remark 3.11 (Nonce Range Rounding and Asymmetry).
When $\gamma_{1}$ is not divisible by $|S|$, rounding slightly reduces the nonce range. This has negligible impact: for ML-DSA-65 with $|S|=17$, the range loss is $2\times 10^{-5}$ per coordinate and does not affect the z-bound check since $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor\geq\gamma_{1}-(|S|-1)>\gamma_{1}-\beta$ (holds for all $|S|\leq\beta=196$, i.e., $|S|<197$, covering all practical configurations). See Appendix G for details.
Additionally, the threshold protocol uses a symmetric per-party nonce range
$\bigl\{-\lfloor\gamma_{1}/|S|\rfloor,\;\ldots,\;\lfloor\gamma_{1}/|S|\rfloor\bigr\},$
whereas FIPS 204 single-signer ML-DSA uses the slightly asymmetric range $\{-\gamma_{1}+1,\ldots,\gamma_{1}\}$ (omitting $\gamma_{1}$). The statistical distance between the symmetric range $\{-\gamma_{1},\ldots,\gamma_{1}\}$ and the FIPS 204 range is $1/(2\gamma_{1}+1)<2^{-20}$ per coordinate, and the full-vector statistical distance over $n\ell=1280$ coefficients is bounded by $1280/(2\gamma_{1}+1)\approx 1.22\times 10^{-3}<2^{-9}$. This difference is negligible and is absorbed into the $\epsilon_{\mathsf{IH}}$ term of Theorem 4.2; it does not affect the security argument or FIPS 204 compatibility of the output signature format.